Monday, September 1, 2014
Our Los Angeles network is dropping packets. We are in contact with the DC. Update: 8:52: The connection is stable again. Update: 20:55: The issue is back again. Update: 21:04: It looks resolved again.
Saturday, August 30, 2014
We have an incoming DDoS attack on dalpure1. It should be taken care of shortly. Update 15:54: This has been solved.
Sunday, August 17, 2014
We are experiencing some packet loss in this network. We are in contact with the DC. Please bare with us.
Friday, August 15, 2014
There is a network issue with our IP subnets in this network zone. We are in contact with the DC. Update: It looks like we are getting affected by this issue: http://blogs.cisco.com/sp/global-internet-routing-table-reaches-512k-milestone/ and getting filtered by some of the internet carriers. Update 23:08: The issue is temporarily fixed. You may experience slowness until tomorrow.
Monday, August 11, 2014
Date: 11 August 2014, Monday Start Time: 2:00 PM PST End Time: 2:30 PM PST Cause: Expansion Expected Downtime: 30 Minutes In order for us to deploy new nodes in our Los Angeles network, we are going to have to move a few servers off their current rack to a new rack. We apologize for the inconvenience this maintenance will cause. Update 1:55 PM (PST): Maintenance has...
Thursday, August 7, 2014
node2 Dallas just kernel panic'ed, we are rebooting it. Update 20:10: All containers have booted. They'll go down for a few moments and come back online while the server recalculates the quotas used. Update 20:40: Everything is back to normal.
Friday, July 18, 2014
We are aware that Los Angeles nodes are down. There is a network issue which the DC is working on. Update 05:30: The network is back up.
Sunday, June 29, 2014
From the announcement we sent out last week: You are receiving this email because you have active offloaded MySQL service. We are writing to inform you of a maintenance window regarding our Dallas offloaded MySQL service. This window is scheduled for Sunday, June 29th from 10 AM to 11 AM CDT (-5 UTC). We do not expect to have a total outage of the service, but connectivity may be...
Tuesday, June 24, 2014
Rebooting caused a kernel panic on this node2 Dallas and we are running fsck at the moment. Please be patient. Update 19:16: fsck is done. The node is booted. VPSes are booting at the moment. To not to wait, go to your client area and boot your VPS.
Tuesday, June 24, 2014
lapure1 is not coming back online. We are working on it. Your patience is appreciated. Update: This node is back online.